# Leveraging RAG-LLMs for Urban Mobility Simulation and Analysis

### v1

@@ I Introduction
Existing commercial e-mobility platforms [3, 4] are already operational across the globe. While these tools are widely used, their proprietary nature imposes significant limitations for researchers and practitioners aiming to explore diverse scenarios and customize solutions. Simulation-based solutions have emerged as a preferable alternative, offering flexibility and adaptability for studying various e-mobility scenarios. In this context, traditional traffic simulation platforms such as AIMSUN [5] and VISSIM [6] primarily address traffic congestion and scheduling, whereas emerging e-mobility solutions focus on energy management and sustainability. These platforms offer some space for customization but are still not open-sourced. Among these, the open-source tool Simulation of Urban MObility (SUMO) has gained widespread adoption in the research community due to its versatility and extensibility [7]. However, despite its strengths, SUMO has certain limitations, particularly in handling large-scale scenarios and integrating with modern cloud-based infrastructures. These limitations become critical in shared e-mobility, where factors such as poor service integration, limited scalability, lack of user-centric solutions, and inadequate energy consumption forecasting remain insufficiently addressed [8].
@@
To address these challenges, there is a growing need for cloud-based SUMO solutions that leverage the scalability of cloud platforms. Recent advancements in cloud-based SUMO focus on enhancing large-scale simulations and real-time data processing, but most studies concentrate on vehicular network optimization rather than shared e-mobility. [9] offers a high-level cloud-based SUMO framework may be limited to simple queries or pure dashboard-based designs which is non-user-friendly. Meanwhile, the emergence of Large Language Model (LLM)s has opened new opportunities for enhancing intelligent decision-making in traffic-related domains, including delivery scheduling, location prediction, and secure communications. Building on this potential, we propose an LLM-powered, cloud-based platform that integrates dynamic traffic simulation with natural language interaction, aiming to bridge the gap between traditional simulation tools and user-friendly, adaptive shared mobility solutions. In particular, our platform supports scenario-based deployment and evaluation of eHub distributions, addressing the lack of scalable simulation toolsets for shared mobility infrastructures. This work builds upon our previous research on energy modeling [10], optimization algorithms [11], and the multi-modal optimization framework for shared e-mobility [12]. Thanks to the modular and data-centric architecture, our platform can be easily adapted for digital twin modeling and deployment in different regions. It separates the system logic from specific datasets, meaning that with appropriate traffic and mapping information, the platform can recreate realistic local mobility patterns without extensive reengineering. The platform’s key contributions are outlined next:
@@
  • Comprehensive Cloud-based Open-Source Shared e-mobility Platform: We present an open-source framework for shared e-mobility with modular components for traffic simulation, docking station deployment, multi-modal transport, energy-aware route planning, and user interaction via an Android app using an agent-in-the-loop approach. A Dockerized architecture ensures cloud scalability. Source code will be released upon paper acceptance.
@@
  • LLM-powered Traffic Simulation Data and Travel History Data Query Enhancement: We present the first Retrieval-Augmented Generation (RAG)-based approach tailored for post-simulation querying over structured traffic simulation data and user travel data. By integrating LLMs with schema-level RAG, our platform enables natural language interaction with simulation outputs for different types of users, including end users and system operators. The schema-level RAG framework supports context-aware Text-to-SQL generation, achieving an average execution accuracy of 0.98 on user queries and 0.81 on system operator queries. This empowers both users and system operators to explore personalized, fine-grained insights going beyond predefined charts or views, and makes simulation data more understandable and accessible in real-world scenarios.
@@
The remainder of the paper is organized as follows: Section II presents a literature review of relevant platforms and provides a comparison with ours. Section III introduces the architecture and details of the proposed platform. introduces the technical details for LLM layer. Section V outlines the experimental setup and analyzes the results. Finally, Section VI discusses the conclusions and future work.
@@ II-B LLM-powered Transportation System
A notable starting point of LLM application in transportation is the work presented in [20], which discusses the utilization of LLMs and RAG in Smart Urban Mobility. However, the study primarily focuses on system design without delving into the implementation details. For traffic scenario generation, the RealGen framework introduced in [21] offers a promising approach. RealGen synthesizes new scenarios by combining behaviors from multiple retrieved examples in a gradient-free manner. These behaviors can originate from predefined templates or tagged scenarios, providing flexibility and adaptability in scenario creation. In the domain of intelligent driving assistance, the IDAS system presented in [22] demonstrates an efficient approach to reading and comprehending car manuals for immediate, context-based aid. This system provides new insights into enhancing driving assistance by leveraging the capabilities of RAG models. Compared to these works, ChatSUMO [23] provides a more comprehensive and operational framework by integrating LLMs into the full pipeline of SUMO traffic scenario generation, and customization, significantly reducing the technical barriers for users without traffic simulation expertise. In contrast, our approach focuses specifically on the post-simulation stage, extending LLM integration on the user side to enable schema-level interaction with structured simulation outputs. While prior work primarily emphasizes data generation or simulation modeling, we target the downstream task of querying and interpreting simulation results. Unlike conventional systems that only support fixed dashboard metrics or static visualizations, our platform uniquely integrates RAG-LLM interfaces to allow flexible, natural language querying over arbitrary simulation results.
@@ III Cloud-Based Platform Architecture
In this section, we present a detailed of the system’s architecture and implementation, focusing on its core components and their interconnections.
@@ III-B3 Dashboard
The interactive dashboard is served as the central interface for controlling and visualizing various aspects of the shared e-mobility emulation platform. It displays real-time traffic simulations for the Dublin City Centre (DCC) map, utilizing GCB as the backend for handling large-scale, time-series traffic data generated by the SUMO system. It ensures high performance and scalability, offering users a seamless way to set up scenarios, control simulations, and analyze traffic patterns in real-time, as illustrated in Fig. 2.
@@ V-A1 Simulation Setup
The DCC map was selected for funding requirements and its prior use in a study on autonomous vehicles. Covering a 5 km x 3.5 km area, the existing SUMO map provided a solid foundation for simulating a typical 24-hour workday using traffic data from Dublin’s SCATS system, which collects vehicle counts every 6 minutes from 480 locations. Data from Transport Infrastructure Ireland, averaged over several months, modeled realistic traffic behavior [24]. To support multi-modal simulations, modifications were made to include pedestrian and bicycle lanes using SUMO’s Netconvert tool. Traffic data from the 2022 Irish census [25], showing 87,000 walkers and 37,000 cyclists, were incorporated, with peak times between 5:30 am and 7 pm.
@@ V-A2 Optimization Module Evaluation Setup
Figure 5: Error Type Proportions Across Different Models and Users
@@
V-A3 RAG Module Evaluation Setup
Figure 6: Comparison of Median Value of Evaluation Metrics for Models on System Operator Queries
Figure 7: Comparison of Median Value of Evaluation Metrics for Models on User Queries
V-B Evaluation and Analysis
V-B1 Optimization Module
The key observations from the results are as follows:
V-B2 RAG module
@@ VI Conclusion and Future Works
This work presents a comprehensive shared e-mobility platform that integrates cloud-based simulation, real-time data processing, and RAG-enhanced decision-making to address challenges in sustainable urban transportation. The platform supports dynamic multi-modal routing and flexible docking-based systems through a scalable and user-centric design. Schema-level RAG evaluation highlights XiYanSQL’s high execution accuracy of 0.81 on system operator queries and 0.98 on user queries. Error analysis reveals that logic errors are the predominant failure type for most models, while granularity issues are more common in user-level queries. It is important to note that, although the LLM-based RAG framework perform robustly overall, generated responses are not always perfect. Users and system operators should treat LLM outputs as supportive decision-making aids rather than definitive conclusions. Based on our evaluation, we recommend cross-checking critical decisions against simulation data or dashboard outputs when high reliability is required. Future work will focus on extending the system to support multi-objective optimization and multi-user scenarios, enabling greater adaptability to diverse real-world demands. As urban mobility becomes increasingly complex and interconnected, integrating public transportation networks into the platform is a natural next step. Additionally, the RAG framework will be enhanced by incorporating external multi-modal knowledge sources to further improve the faithfulness and contextual grounding of generated responses. Beyond these directions, we also plan to extend the platform toward digital twin modeling by linking to real-world transportation data and live traffic conditions. While the current platform focuses on arbitrary scenario generation with predefined traffic flow configurations using historical traffic data, future enhancements will aim to support more dynamic and real-world simulations and multi-modal data inputs, including text and images, to further improve realism and decision-making support.

### v2

@@ I Introduction
Nowadays, commercial e-mobility platforms such as Bird and Moovit have been widely deployed [3]. However, their proprietary nature imposes significant limitations on researchers and practitioners aiming to explore diverse scenarios and customize solutions. Simulation-based solutions have emerged as a preferable alternative, offering flexibility and adaptability for studying various e-mobility scenarios. In this context, traditional traffic simulation platforms such as AIMSUN [4] and VISSIM [5] primarily address traffic congestion and scheduling, whereas emerging e-mobility solutions focus on energy management and sustainability. These platforms offer some scope for customization but are still not open-sourced. Among these, the open-source tool Simulation of Urban MObility (SUMO) has gained widespread adoption in the research community due to its versatility and extensibility [6]. However, despite its strengths, SUMO has certain limitations, particularly in handling large-scale scenarios and integrating with modern cloud-based infrastructures. These limitations become critical in shared e-mobility, where factors such as poor service integration, limited scalability, lack of user-centric solutions, and inadequate energy consumption forecasting remain insufficiently addressed [7].
@@
To address these challenges, there is a growing need for cloud-based SUMO solutions that leverage the scalability of cloud platforms. Recent advancements in cloud-based SUMO focus on enhancing large-scale simulations and real-time data processing, but most studies concentrate on vehicular network optimization rather than shared e-mobility. [8] offers a high-level cloud-based SUMO framework, but it may be limited to simple queries or purely dashboard-based designs, which are not user-friendly. Meanwhile, the emergence of Large Language Models (LLMs) has opened new opportunities for enhancing intelligent decision-making in traffic-related domains, including delivery scheduling, location prediction, and secure communications. Building on this potential, we propose an LLM-powered, cloud-based platform that integrates dynamic traffic simulation with natural language interaction, aiming to bridge the gap between traditional simulation tools and user-friendly, adaptive shared mobility solutions. In particular, our platform supports scenario-based deployment and evaluation of eHub distributions, addressing the lack of scalable simulation toolsets for shared mobility infrastructures. This work builds upon our previous research on energy modeling [9, 10], optimization algorithms [3], and the multi-modal optimization framework for shared e-mobility [11]. Thanks to the modular and data-centric architecture, our platform can be easily adapted for digital twin modeling and deployment in different regions. It separates the system logic from specific datasets, meaning that with appropriate traffic and mapping information, the platform can recreate realistic local mobility patterns without extensive reengineering. The platform’s key contributions are outlined next:
@@
  • A comprehensive, open-source, cloud-based platform for shared e-mobility: We present a framework for shared e-mobility with modular components for traffic simulation, docking station deployment, multi-modal transport, energy-aware route planning, and user interaction via an Android app using an agent-in-the-loop approach. A Dockerized architecture ensures cloud scalability. Source code is available  [^1] https://github.com/SFIEssential/Essential.
@@
  • Enhanced querying of simulation and travel data through LLMs: We present the first Retrieval-Augmented Generation (RAG)-based approach tailored for post-simulation querying over structured traffic simulation data and user travel data. By integrating LLMs with schema-level RAG, our platform enables natural language interaction with simulation outputs for different types of users, including end users and system operators. The schema-level RAG framework supports context-aware Text-to-SQL generation, achieving an average execution accuracy of 0.98 on user queries and 0.81 on system operator queries. This empowers both users and system operators to explore personalized, fine-grained insights going beyond predefined charts or views, and makes simulation data more understandable and accessible in real-world scenarios.
@@
The rest of the paper is structured as follows: Section II reviews related platforms and compares them with ours. Section III describes the architecture of the proposed platform. Section IV details the implementation of the LLM layer. Section V presents the experimental setup and analyzes the results. Finally, Section VI concludes the paper and outlines directions for future work.
@@ II-B LLM-powered Transportation System
A notable starting point for applying LLMs in transportation is presented in [19], which discusses the utilization of LLMs and RAG in Smart Urban Mobility. However, the study primarily focuses on system design without delving into the implementation details. For traffic scenario generation, the RealGen framework introduced in [20] offers a promising approach. RealGen synthesizes new scenarios by combining behaviors from multiple retrieved examples in a gradient-free manner. These behaviors can originate from predefined templates or tagged scenarios, providing flexibility and adaptability in scenario creation. In the domain of intelligent driving assistance, the IDAS system presented in [21] demonstrates an efficient approach to reading and comprehending car manuals for immediate, context-based aid. This system provides new insights into enhancing driving assistance by leveraging the capabilities of RAG models. Compared to these works, ChatSUMO [22] provides a more comprehensive and operational framework by integrating LLMs into the full pipeline of SUMO traffic scenario generation, and customization, significantly reducing the technical barriers for users without traffic simulation expertise. In contrast, our approach focuses specifically on the post-simulation stage, extending LLM integration on the user side to enable schema-level interaction with structured simulation outputs. While prior work primarily emphasizes data generation or simulation modeling, we target the downstream task of querying and interpreting simulation results. Unlike conventional systems that only support fixed dashboard metrics or static visualizations, our platform uniquely integrates RAG-LLM interfaces to allow flexible, natural language querying over arbitrary simulation results.
@@ III Cloud-Based Platform Architecture
In this section, we present a detailed overview of the system’s architecture and implementation, focusing on its core components and their interconnections.
@@ III-B3 Dashboard
The interactive dashboard serves as the central interface for controlling and visualizing various aspects of the shared e-mobility emulation platform. It displays real-time traffic simulations for the Dublin City Centre (DCC) map, utilizing GCB as the backend for handling large-scale, time-series traffic data generated by the SUMO system. It ensures high performance and scalability, offering users a seamless way to set up scenarios, control simulations, and analyze traffic patterns in real-time, as illustrated in Fig. 2.
@@ V-A1 Simulation Setup
Our simulation was based on the DCC map, chosen for its links to prior research in this field [23]. The map, covering a 5 km x 3.5 km area, provided a robust foundation for simulating a 24-hour workday, using vehicle counts collected every 6 minutes from 480 locations via Dublin’s SCATS system. Data from Transport Infrastructure Ireland, averaged over several months, modeled realistic traffic behavior. To support multi-modal simulations, modifications were made to include pedestrian and bicycle lanes using SUMO’s Netconvert tool. Traffic data from the 2022 Irish census [24], showing 87,000 walkers and 37,000 cyclists, were incorporated, with peak times between 5:30 am and 7 pm.
@@ V-B Evaluation and Analysis
Figure 5: Comparison of Median Value of Evaluation Metrics for Models on System Operator Queries.
Figure 6: Comparison of Median Value of Evaluation Metrics for Models on User Queries.
V-B1 Optimization Module
The key observations from the results are as follows:
V-B2 RAG module
Figure 7: Error Type Proportions Across Different Models and Users.
@@ VI Conclusion and Future Works
This paper presents a comprehensive shared e-mobility platform that integrates cloud-based simulation, real-time data processing, and RAG-enhanced decision-making to address the challenges in sustainable urban transportation. The platform supports dynamic multi-modal routing and flexible docking-based systems through a scalable and user-centric design. The schema-level RAG evaluation highlights XiYanSQL’s high execution accuracy of 0.81 on system operator queries and 0.98 on user queries. An error analysis reveals that logic errors are the predominant failure type for most models, while granularity issues are more common in user-level queries. Although the LLM-based RAG framework performs robustly, the generated responses are not always perfect. Users and system operators should treat LLM outputs as supportive decision-making aids rather than definitive conclusions and should cross-check critical decisions against simulation data or dashboard outputs when high reliability is required.
Finally, some limitations remain to be addressed in this work. First, the platform’s scalability under high user loads and real-time traffic updates has not been fully assessed. Second, its adaptability to different urban networks and traffic configurations requires further investigation. Third, usability testing has yet to be conducted. As part of the future work, we will address these gaps by focusing on large-scale performance, cross-platform portability, integration with public transportation, and usability studies with a view to further enhance user interaction and system responsiveness.